OpenStudy (anonymous):

A piece of wire 12 m long is cut into two pieces. One piece is bent into the shape of a circle of radius r and the other is bent into a square of side s. How should the wire be cut so that the total area enclosed is: a) a maximum r= and s= ? b) a minimum r= and s= ?

OpenStudy (anonymous):

hero got it

hero (hero):

Thanks for having so much confidence in me

OpenStudy (anonymous):

no calc for this one, pure reason gets it i think

OpenStudy (anonymous):

think square

OpenStudy (anonymous):

s=\[\pi r ^{2}+(12-r)^{2}=\S\] for the max u should ds/dr =0 so u could find r

OpenStudy (anonymous):

sri u can't use 12-r as jero said 2pi r + s=12 , u should find r from this equation

OpenStudy (anonymous):

sri hero ! i make a mistake when i typed u name

hero (hero):

Don't pay any attention to the equations I wrote

hero (hero):

Actually, that equation is wrong

OpenStudy (anonymous):

maximum, make a circle. minimum, make a square

OpenStudy (asnaseer):

for a given length of wire \(l\), if it is bent into a circle we get:\[2\pi r=l\quad\implies \pi r=\frac{l}{2}\quad\implies Area=\pi r^2=\frac{l^2}{4}\]and if bent into a square with side length \(s\) we get:\[4s=l\quad\implies s=\frac{l}{4}\quad\implies Area=s^2=\frac{l^2}{16}\]so, as @satellite73 said, maximum area is achieved for a circle, minimum for a square.

hero (hero):

You should explain the minimum and maximum part

OpenStudy (asnaseer):

sorry - mistake in area of circle, should be:\[2\pi r=l\quad\implies r=\frac{l}{2\pi}\quad\implies Area=\pi r^2=\frac{l^2}{4\pi}\]

OpenStudy (asnaseer):

and:\[\frac{l^2}{4\pi}>\frac{l^2}{16}\]since \(\pi\) is less than 4.

hero (hero):

You kind of short-cutted the explanation by using l instead of 12 - s

OpenStudy (asnaseer):

I was trying to explain that, for a given length of wire, the circle will enclose a bigger area than a square.

hero (hero):

The length for the circle is different from the length for the square, but yet you use the same variable for both without distinguishing between the two

OpenStudy (asnaseer):

for a given length of wire, if we bend it into a circle we will get a larger area than if we bend it into a square. so the next steps would be to cut the wire of length 12 into two such that one piece is of length x and the other is of length 12-x and work out what gives us a maximum/minimum enclosed area. what I showed was just the first step to establish that a circle always encloses a larger area than a square for a given length of wire.

OpenStudy (asnaseer):

therefore maximum is achieved when entire wire is used to create a circle where:\[2\pi r=12\quad\implies r=\frac{6}{\pi}\text{ and }s=0\]and minimum area is achieved when the entire wire is used to create a square where:\[4s=12\quad\implies s=3\text{ and }r=0\]

hero (hero):

You kinda skipped a whole bunch of steps there, but okay

OpenStudy (anonymous):

I entered the minimum but it was wrong...

hero (hero):

What minimum did you enter?

OpenStudy (anonymous):

s=0 r =3?

OpenStudy (anonymous):

wait i mean the other way..

OpenStudy (anonymous):

s= 3 and r=0...

hero (hero):

for a) it refers to a maximum and we know the circle has the maximum therefore, so s = 12 r = 0 for b) it refers to a minimum and we know that the square has a minimum s = 0, but the square cannot be defined in terms of r because it doesn't have a radius.

OpenStudy (asnaseer):

ok guys - looks like this problem was a bit more subtle than I thought. the problem states find the maximum and minimum "total area enclosed". so if asume we cut the 12m wire into two with one wire of length \(x\) used for the square and the remaining wire of length \(12-x\) used for the circle, then we get:\[2\pi r=12-x\quad\implies r=\frac{12-x}{2\pi}\quad\implies A_{circle}=\pi r^2=\frac{(12-x)^2}{4\pi}\]and:\[4s=x\quad\implies s=\frac{x}{4}\quad\implies A_{square}=s^2=\frac{x^2}{16}\]so the total area A is given by:\[A=A_{circle}+A_{square}=\frac{(12-x)^2}{4\pi}+\frac{x^2}{16}\]to get the minimum/maximum we need to set the differential of A to zero, so:\[A'=\frac{-2(12-x)}{4\pi}+\frac{2x}{16}=\frac{x}{8}-\frac{12-x}{2\pi}=0\]this gives:\[\frac{x}{8}=\frac{12-x}{2\pi}\]\[2\pi x=96-8x\]\[\pi x=48-4x\]\[(4+\pi)x=48\]\[x=\frac{48}{4+\pi}\]the second derivative of A gives us:\[A''=\frac{1}{8}+\frac{1}{2\pi}>0\]therefore we know \(x=\frac{48}{4+\pi}\) MUST represent a minimum value. we therefore get the following for the minimum:\[x=\frac{48}{4+\pi}\]\[r=\frac{12-x}{2\pi}=\frac{12-\frac{48}{4+\pi}}{2\pi}=\frac{\frac{48+12\pi-48}{4+\pi}}{2\pi}=\frac{12\pi}{2\pi(4+\pi)}=\frac{6}{4+\pi}\]\[s=\frac{x}{4}=\frac{12}{4+\pi}\] the maximum remains at using the entire wire for the circle so \(r=\frac{6}{\pi}\), \(s=0\).

OpenStudy (anonymous):

thanks! ^-^

OpenStudy (asnaseer):

yw